# Bayes, E-values, and Testing

### v3

@@ 3 E-values and E-processes
Remark 3.1 (Representation versus construction).
Two layers should be kept separate: (i) the representation layer, where coherent prediction under log-loss identifies the likelihood ratio as the canonical evidence statistic, with boundary selection governed by moderate deviations of $\log B_{t}$; and (ii) the construction layer, where any nonnegative supermartingale with $\operatorname{\mathbb{E}}_{0}[E_{0}]\leq 1$ is a valid E-process, with Markov/Ville certifying Type I control.
@@ 3.1 E-variables
Unlike $p$-values, E-values admit natural combination rules: independent E-variables can be multiplied, and possibly dependent E-variables can be averaged, both preserving the E-variable property [41, 21].
@@ 3.2 E-processes and optional stopping
Thus a sequential test rejecting $H_{0}$ at the first time $E_{t}\geq 1/\alpha$ controls Type I error at level $\alpha$ under arbitrary optional stopping [21].
Validity versus optimality.
Markov’s inequality yields the anytime-valid rejection rule $\tau=\inf\{t:E_{t}\geq 1/\alpha\}$, guaranteeing Type I control, but does not by itself select a Bayes-optimal rejection region. Optimality enters through the decision problem: the cutoff is $c^{\star}=\pi_{0}L_{10}/(\pi_{1}L_{01})$; see Section˜3.7.
@@ 3.3 Canonical construction: likelihood ratio E-process
$\operatorname{\mathbb{E}}_{H_{0}}[e_{s}\,|\,\mathcal{F}_{s-1}]=\sum_{x_{s}}\frac{q(x_{s}\,|\,X^{s-1})}{p_{0}(x_{s}\,|\,X^{s-1})}\cdot p_{0}(x_{s}\,|\,X^{s-1})=\sum_{x_{s}}q(x_{s}\,|\,X^{s-1})=1.$
@@
Validity-only constructions satisfying $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ can yield genuine E-processes without arising from predictive kernels [33, 31], but they do not inherit the Bayes-risk semantics that the log-loss bridge provides (Section˜2.3). Cutoffs belong to the decision layer (Section˜3.7), not the evidence layer.
@@ 3.4 Mixture E-processes: the Bayesian construction
Proposition 3.9 (Bayes factor is an E-variable).
@@
satisfies $\operatorname{\mathbb{E}}_{M_{0}}[\text{BF}_{10}]=1$, and is therefore an E-variable under $M_{0}$.
@@
Hence sequential Bayes factors inherit anytime validity, multiplicative combination, and Ville’s inequality.
@@ 3.5 Worked example: Bernoulli coin with Beta prior
3.6 Good’s paradox and the stopping rule principle
The stopping-rule dependence of classical procedures, and independence of Bayesian and E-value methods, was central in debates. By the law of the iterated logarithm, a frequentist using optional stopping can always reject a true null given patience [18].
@@ 3.6 Good’s paradox and the stopping rule principle
E-values clarify this tension: since Bayes factors are E-values (Proposition˜3.9) and E-processes maintain Type I error control under optional stopping (Theorem˜3.5), one obtains both likelihood-based evidence and sequential validity without pre-specifying stopping rules. The “safe testing” programme formalizes this perspective [20, 38].
@@ 3.8 Critical region geometry for E-values
The critical region for sequential evidence is governed by two distinct asymptotic geometries. Conflating them is the source of the conservatism noted in Remark˜2.7.
@@
Tail scale Polynomial: $\leq 1/y$ Exponential: $\sqrt{(\log y)/t}$
@@
$\operatorname{\mathbb{P}}_{0}\!\left(\sup_{s\leq t}E_{s}\geq y\right)\leq\frac{1}{y}.$
Markov/Ville certify validity for a chosen cutoff; setting $y=1/\alpha$ gives the anytime-valid certificate $\operatorname{\mathbb{P}}_{0}(\sup E_{s}\geq 1/\alpha)\leq\alpha$. This bound operates on $E_{t}$ at a polynomial $\sim\!1/y$ scale. It guarantees Type I control but is agnostic about likelihood-ratio structure and therefore does not optimize the critical region.
@@
and moderate-deviation theory governs the fluctuations: deviations of $(1/t)\log B_{t}$ from its KL slope are of order $\sqrt{(\log y)/t}$—an exponentially sharper scale than the polynomial $1/y$ of Markov/Ville. This is the appropriate scaling for boundary selection and power analysis. [^2] A concrete comparison: at $t=100$ observations and threshold $y=20$, the Markov/Ville bound gives $\operatorname{\mathbb{P}}_{0}(E_{t}\geq 20)\leq 1/20=0.05$, while the moderate-deviation fluctuation scale is $\sqrt{(\log 20)/100}\approx 0.17$ nats—the two operate on entirely different objects ($E_{t}$ versus $(1/t)\log B_{t}$) and at different scales. See also Proposition˜5.4.
@@
The Markov/Ville boundary is polynomial and conservative; the likelihood-ratio boundary is governed by moderate-deviation theory and yields exponentially sharper boundaries. Procedures relying solely on Markov/Ville calibration may therefore be conservative in finite samples, since they do not exploit the KL growth structure available for likelihood-ratio processes. The evidence process $B_{t}$ is the same in both cases; what differs is the geometry used to select and evaluate the critical region.
@@
Rate statements are naturally expressed on the log scale: we study $(1/t)\log E_{t}$ rather than $E_{t}/t$.
@@
The $\sqrt{(\log y)/t}$ moderate-deviation scale is not merely descriptive: in related Bayes-risk calibrations one can solve explicitly for the optimizing critical level, yielding thresholds of order $\sqrt{(\lambda+m+1)\log n}$ rather than $1/n$-type calibrations, where $\lambda,m$ are problem-specific parameters; see Datta et al. [9] for a worked-out instance.
@@ 3.10 Comparison with $p$-values
Ville’s inequality provides valid Type I error control at arbitrary stopping times; fixed-horizon $p$-values require adjustments (e.g., $\alpha$-spending) for sequential use.
@@
Independent E-variables can be multiplied; dependent ones can be averaged. Combining $p$-values requires distributional assumptions or conservative bounds (Bonferroni, Fisher).
Relation to Bayes factors.
Bayes factors are E-variables under the null (Proposition˜3.9), inheriting all sequential validity properties [20].
@@ 3.11 Duality with confidence sequences
Bayes factors arise as likelihood-ratio E-processes after marginalization over parameters, retaining anytime-valid error control provided the construction is predictably normalized.
4 Composition, amalgamation, and obstructions in the evidence cone
This section formalizes the algebraic structure of the evidence cone: which operations on E-processes preserve validity, and which provably do not.
4.1 The evidence cone and its basic algebra
Definition 4.1 (Evidence cone under $H_{0}$).
Let $H_{0}$ denote the null model class (as in §3, Definition˜3.2). Define the evidence cone
@@ 4.1 The evidence cone and its basic algebra
Similarly, on a filtered space $(\Omega,\mathcal{F},(\mathcal{F}_{t}),\operatorname{\mathbb{P}})$, define the process evidence cone
@@
Proposition˜4.2 shows that validity of averaging and mixtures is a closure property of the evidence cone. Recent work on combining and merging E-values [41, 33] studies finer admissibility questions, which merging functions are optimal within the cone, but the basic algebra captured here underlies all such results.
@@ 4.2 Anytime validity as a compositional consequence of Ville
Remark 4.5 (Anytime validity).
“Anytime-valid” inference is the statement of Theorem˜4.4, following from nonnegative supermartingale structure and Ville’s inequality. Likelihood-ratio and Bayes-factor processes are anytime-valid by construction (Propositions˜3.6 and 3.9).
@@ 4.3 Sequential composition (stitching) of evidence processes
Under Definition˜4.6, $(\tilde{E}_{t})_{t\geq 0}\in\mathcal{E}^{\mathrm{proc}}(H_{0})$. Thus $\tilde{E}_{\tau}$ and $\sup_{t}\tilde{E}_{t}$ satisfy the Ville bounds of Theorem˜4.4.

### v4

@@ 3 E-values and E-processes
Remark 3.1 (Representation versus construction—a guiding distinction).
Throughout this paper, the reader should keep two layers separate. Representation layer: under coherent prediction and log-loss, the Fubini/Bayes-risk decomposition identifies the likelihood ratio (or Bayes factor) as the canonical evidence statistic; boundary selection is governed by the moderate-deviation geometry of $\log B_{t}$ (scale $\sqrt{(\log b)/t}$). Construction layer: any nonnegative supermartingale with $\operatorname{\mathbb{E}}_{0}[E_{0}]\leq 1$ is a valid E-process; Markov/Ville calibration certifies Type I control (scale $1/b$) regardless of likelihood-ratio structure. Validity (Markov/Ville) and optimality (LR/Bayes risk) are complementary, not interchangeable.
@@ 3.1 E-variables
Unlike $p$-values, E-values admit natural combination rules. If $E_{1},\dots,E_{m}$ are independent E-variables for the same null, their product $\prod_{i=1}^{m}E_{i}$ is again an E-variable (by the product property of expectations under independence). If $E_{1},\dots,E_{m}$ are possibly dependent E-variables for the same null, their average $\frac{1}{m}\sum_{i=1}^{m}E_{i}$ is an E-variable (by linearity of expectation). These properties make E-values a flexible currency for combining evidence across experiments and analyses [42]. Recent work has emphasized E-values as a replacement for classical $p$-values in sequential hypothesis testing, highlighting their supermartingale structure and optional-stopping validity [22].
@@ 3.2 E-processes and optional stopping
Ville’s inequality is the foundation for anytime-valid inference: a sequential test that rejects $H_{0}$ at the first time $E_{t}\geq 1/\alpha$ controls Type I error at level $\alpha$ under arbitrary optional stopping, without any pre-specification of the sample size or stopping rule. This supermartingale perspective underlies recent proposals to reformulate hypothesis testing in terms of E-values rather than fixed significance levels [22].
Validity is not optimality.
Markov’s inequality yields $\operatorname{\mathbb{P}}_{0}(E\geq c)\leq 1/c$, hence the anytime-valid rejection rule $\tau=\inf\{t:E_{t}\geq 1/\alpha\}$. Markov/Ville inequalities turn an E-process into an anytime-valid rejection rule, guaranteeing Type I control, but they do not by themselves select a Bayes-optimal rejection region. Optimality enters through the decision problem (priors and losses), in which the likelihood ratio is the canonical statistic and the cutoff is $c^{\star}=\pi_{0}L_{10}/(\pi_{1}L_{01})$; see Section˜3.7. The rule “reject when $E_{t}\geq 1/\alpha$” is the risk-budget special case of this decision principle.
@@ 3.3 Canonical construction: likelihood ratio E-process
$\operatorname{\mathbb{E}}_{H_{0}}[e_{s}\,|\,\mathcal{F}_{s-1}]=\sum_{x_{s}}\frac{q(x_{s}\,|\,X^{s-1})}{p_{0}(x_{s}\,|\,X^{s-1})}\cdot p_{0}(x_{s}\,|\,X^{s-1})\\ =\sum_{x_{s}}q(x_{s}\,|\,X^{s-1})=1.$
@@
The condition $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ (and its process analogue) is a validity constraint: it certifies anytime Type I control via Markov/Ville. A distinct issue is representation: when evidence is required to arise from a coherent predictive law under log-loss, the additive evidence is a log likelihood ratio (Section˜2.3), and the canonical E-process is the corresponding likelihood-ratio martingale. Validity-only constructions that satisfy $\operatorname{\mathbb{E}}_{0}[E]\leq 1$ but do not arise from predictive kernels can yield genuine E-processes—in particular, for composite nulls where no single likelihood ratio suffices [1, 33]—but they do not inherit the Bayes-risk/weight-of-evidence semantics that the log-loss bridge provides. Cutoffs, in turn, belong to the decision layer (Section˜3.7), not the evidence layer.
@@ 3.4 Mixture E-processes: the Bayesian construction
Proposition 3.9 (Bayes factor is an E-variable under $M_{0}$).
@@
satisfies $\operatorname{\mathbb{E}}_{M_{0}}[\text{BF}_{10}]=1$, and is therefore an E-variable under the prior-predictive (mixture) null model $M_{0}$.
@@
This establishes validity under the model-averaged null $M_{0}$; for composite null classes $\mathcal{P}$, uniform validity $\sup_{P\in\mathcal{P}}\operatorname{\mathbb{E}}_{P}[E_{t}]\leq 1$ is a strictly stronger requirement and generally distinct from mixture-null validity. Hence sequential Bayes factors inherit anytime validity, multiplicative combination, and Ville’s inequality.
Composite nulls: mixture validity vs uniform validity.
When the null is specified as a set of measures $\mathcal{P}$ (a composite hypothesis), the e-value literature typically requires uniform validity, $\operatorname{\mathbb{E}}_{P}[E_{t}]\leq 1$ for every $P\in\mathcal{P}$, which is strictly stronger than validity under a prior-predictive mixture null $M_{0}=\int P_{\theta}\,\pi_{0}(d\theta)$ (the setting of Proposition˜3.9). In this uniform regime, generic Bayes factors need not be e-variables for $\mathcal{P}$. Nevertheless, a likelihood-ratio representation remains canonical in an appropriate minimax sense: Larsson et al. [27] show that for arbitrary $\mathcal{P}$ and fixed alternative $Q$, there exists a log-optimal (“numeraire”) e-variable $X^{*}$ that can be represented as a likelihood ratio $dQ/dP^{*}$ for a distinguished representative $P^{*}$ (a generalized reverse information projection/bipolar construction). This complements our mixture-null (Bayes-risk) representation: both regimes yield LR-type canonical objects, but with different null semantics. Accordingly, statements of the form “BF == e-value” are correct under the mixture (prior-predictive) null semantics, and generally false under uniform composite-null semantics unless one uses a numeraire/projection construction (or very special priors).
@@ 3.5 Worked example: Bernoulli coin with Beta prior
3.6 Good’s paradox for p-values; stopping-rule invariance via E-processes
Good highlighted a paradox for p-values under optional stopping: the stopping-rule dependence of classical procedures, and independence of Bayesian and E-value methods, was central in debates. By the law of the iterated logarithm, a frequentist using optional stopping can always reject a true null given patience [19].
@@ 3.6 Good’s paradox for p-values; stopping-rule invariance via E-processes
E-values clarify this tension by providing frequentist error control that is valid under arbitrary stopping while also coinciding with Bayesian evidence. Since Bayes factors are E-values under the mixture (prior-predictive) null (Proposition˜3.9), and E-processes maintain Type I error control under optional stopping via Ville’s inequality (Theorem˜3.5), the E-value framework shows that one can have both: evidential measures that are likelihood-based and sequential validity that does not require pre-specification of stopping rules. The “safe testing” programme formalizes this perspective, providing E-value constructions with explicit robustness guarantees [21, 39].
In particular, Good’s example highlights how fixed-horizon $p$-value calibration can break under repeated peeking unless adjusted. E-processes, by contrast, satisfy Ville’s inequality under arbitrary stopping, so Type I error control is preserved regardless of the stopping rule. The remaining question for E-values is not stopping validity but boundary selection: given that validity is guaranteed, which threshold yields Bayes-risk-optimal decisions? Our distinction in the sequel concerns precisely this optimality layer: Markov/Ville certify validity for a chosen boundary, but do not by themselves determine Bayes-risk-aligned rejection regions when LR structure is available.
@@ 3.8 Critical region geometry for E-values
The critical region for sequential evidence is governed by two distinct asymptotic geometries. Failing to distinguish them is the source of the efficiency gap noted in Remark˜2.7.
@@
Tail scale Polynomial: $\leq 1/b$ Exponential: $\sqrt{(\log b)/t}$
@@
Here $b$ is a fixed evidence cutoff (e.g. $b=1/\alpha$); all asymptotics are in the sample size $t$. The $\log b$ term arises solely from the boundary inversion $B_{t}\geq b\Leftrightarrow\log B_{t}\geq\log b$ and is not an independent asymptotic variable.
@@
$\operatorname{\mathbb{P}}_{0}\!\left(\sup_{s\leq t}E_{s}\geq b\right)\leq\frac{1}{b}.$
Markov/Ville certify validity for a chosen cutoff; setting $b=1/\alpha$ gives the anytime-valid certificate $\operatorname{\mathbb{P}}_{0}(\sup E_{s}\geq 1/\alpha)\leq\alpha$. This bound operates on $E_{t}$ at a polynomial $\sim\!1/b$ scale. It guarantees Type I control but is agnostic about likelihood-ratio structure and therefore does not optimize the critical region.
@@
and moderate-deviation theory governs the fluctuations: deviations of $(1/t)\log B_{t}$ from its KL slope are of order $\sqrt{(\log b)/t}$—an exponentially sharper scale than the polynomial $1/b$ of Markov/Ville. The connection is direct: crossing $B_{t}\geq b$ is equivalent to $\log B_{t}\geq\log b$, and since $(1/t)\log B_{t}\approx D_{\operatorname{KL}}(P_{1}\|P_{0})$ under $P_{1}$, the threshold $\log b$ is naturally calibrated against $t\,D_{\operatorname{KL}}$. This is the appropriate scaling for boundary selection and power analysis. [^2] A concrete comparison: for the fixed threshold $b=20$ at $t=100$ observations, the Markov/Ville bound gives $\operatorname{\mathbb{P}}_{0}(E_{t}\geq 20)\leq 1/20=0.05$, while the moderate-deviation fluctuation scale is $\sqrt{(\log 20)/100}\approx 0.17$ nats—the two operate on entirely different objects ($E_{t}$ versus $(1/t)\log B_{t}$) and at different scales. See also Proposition˜5.4.
@@
The Markov/Ville boundary is polynomial and universally valid but may be loose; the likelihood-ratio boundary is governed by moderate-deviation theory and yields exponentially sharper boundaries. Procedures relying solely on Markov/Ville calibration may therefore be potentially loose in finite samples, since they do not exploit the KL growth structure available for likelihood-ratio processes. The evidence process $B_{t}$ is the same in both cases; what differs is the geometry used to select and evaluate the critical region.
In short: Ville/Markov tells you that a boundary $b=1/\alpha$ is valid; it does not tell you it is optimal. Optimal boundaries track $\log B_{t}$ and therefore inherit $\sqrt{(\log b)/t}$-scale moderate deviations around the KL slope.
@@
Rate statements are naturally expressed on the log scale: we study $(1/t)\log E_{t}$ (or $(1/t)\log B_{t}$) rather than $E_{t}/t$. This places the Markov/Ville calibration (on $E_{t}$) and KL growth (on $\log E_{t}$) in a common time-normalized comparison.
Why the Markov region is not an evidential region.
The inequality $\operatorname{\mathbb{P}}_{0}(\sup_{s\leq t}E_{s}\geq 1/\alpha)\leq\alpha$ is a validity certificate for an execution rule; it does not convert $E_{t}$ into a tail probability nor identify a Bayes-optimal rejection region. Interpreting $E_{t}$ as “1/p1/p” conflates two typed objects: Bayes factors compare models relatively via likelihood ratios, whereas Markov/Ville bounds are one-sided guarantees that ignore the alternative and therefore cannot optimize Type II error along the frontier.
@@
The $\sqrt{(\log b)/t}$ moderate-deviation scale is not merely descriptive: in related Bayes-risk calibrations one can solve explicitly for the optimizing critical level, yielding thresholds of order $\sqrt{(\lambda+m+1)\log n}$ rather than $1/n$-type calibrations, where $\lambda,m$ are problem-specific parameters; see Datta et al. [10] for a worked-out instance.
3.9 Validity vs. efficiency: boundary design under LR structure
E-process thresholds such as $E_{\tau}\geq 1/\alpha$ arise from Ville’s inequality and provide exact Type I control under arbitrary stopping. This is a validity-layer guarantee: it does not depend on distributional assumptions beyond the e-process property.
However, when the evidence process admits a likelihood-ratio representation (simple LR, mixture Bayes factor, or numeraire LR against a representative $P^{*}$), the distribution of $\log E_{t}$ under alternatives exhibits a large-deviation and moderate-deviation structure. In that regime, boundary selection becomes an efficiency problem rather than a validity problem.
Specifically, if $\log E_{t}$ satisfies a moderate deviation principle with rate governed by the KL slope, then calibrating thresholds solely via the universal $1/\alpha$ scale may sacrifice efficiency relative to boundaries tuned to the moderate-deviation geometry.
The distinction is therefore not between correct and incorrect thresholds, but between universal validity guarantees and asymptotically efficient boundary design when LR structure is available.
When $E_{t}$ has LR-type structure (simple LR, mixture BF, or numeraire LR against a representative $P^{*}$), the boundary problem reduces to calibrating deviations of $\log E_{t}$, which inherit the KL geometry of the underlying model. Thus the appropriate scaling is dictated by moderate-deviation behavior rather than solely by Markov bounds.
This observation does not contradict the validity of Ville-calibrated e-processes; rather, it identifies conditions under which sharper boundary design is possible.
Validity Efficiency
Ville/Markov Exact Possibly loose
MDP/KL calibration Exact (when e-process) KL-optimal under LR structure
@@ 3.11 Comparison with $p$-values
If $(E_{n})$ is an E-process, Ville’s inequality provides valid Type I error control at arbitrary stopping times. Fixed-horizon $p$-values generally require adjustments (e.g., $\alpha$-spending) for sequential use.
@@
Independent E-variables can be multiplied; possibly dependent E-variables can be averaged. Both operations preserve the E-variable property. Combining $p$-values requires distributional assumptions or conservative bounds (e.g., Bonferroni, Fisher’s method).
Relation to Bayes factors and safe tests.
Bayes factors are E-variables under the mixture (prior-predictive) null (Proposition˜3.9), inheriting all sequential validity properties. The safe testing framework studies E-value constructions with robustness guarantees under optional stopping and model misspecification [21].
@@ 3.12 Duality with confidence sequences
From a Bayesian perspective, Bayes factors arise as likelihood-ratio E-processes after marginalization over parameters. E-values extend this viewpoint by retaining anytime-valid error control under optional stopping while remaining interpretable as sequential likelihood ratios or mixtures thereof, provided the construction is predictably normalized.
Our point in the following sections is not to oppose e-values or the broader e-process programme, but to separate (i) validity certificates (Markov/Ville) from (ii) representational semantics (likelihood-ratio structure under coherent prediction) and (iii) decision-layer boundary choice (Bayes-risk-optimal cutoffs). We view these as complementary, not competing, layers of sequential inference.
4 Composition, amalgamation, and obstructions in the evidence class
This section formalizes the algebraic structure of the evidence class: which operations on E-processes preserve validity, and which provably do not.
4.1 The evidence class and its basic algebra
Definition 4.1 (Evidence class under $H_{0}$).
Let $H_{0}$ denote the null model class (as in §3, Definition˜3.2). Define the evidence class
@@ 4.1 The evidence class and its basic algebra
Similarly, on a filtered space $(\Omega,\mathcal{F},(\mathcal{F}_{t}),\operatorname{\mathbb{P}})$, define the process evidence class
@@
Remark 4.2 (Terminology: class, not cone).
The set $\mathcal{E}(H_{0})$ is convex but not a cone in the strict algebraic sense, since the closure $cE\in\mathcal{E}(H_{0})$ holds only for scalars $c\in(0,1]$; scaling by $c>1$ may violate $\operatorname{\mathbb{E}}_{H_{0}}[cE]\leq 1$. We therefore use “class” rather than “cone.”
@@
Proposition˜4.3 shows that validity of averaging and mixtures is a closure property of the evidence class. Recent work on combining and merging E-values [42, 1] studies finer admissibility questions, which merging functions are optimal within the class, but the basic algebra captured here underlies all such results.
@@ 4.2 Anytime validity as a compositional consequence of Ville
Remark 4.6 (What is “anytime validity”?).
What is called “anytime-valid” or “always-valid” inference is the statement of Theorem˜4.5: it follows from nonnegative supermartingale structure and Ville’s inequality (Theorem˜3.5). In particular, likelihood-ratio and Bayes-factor processes are anytime-valid by construction (Propositions˜3.6 and 3.9), and convex mixtures preserve this property (Proposition˜4.3).
@@ 4.3 Sequential composition (stitching) of evidence processes
Under Definition˜4.7, $(\tilde{E}_{t})_{t\geq 0}\in\mathcal{E}^{\mathrm{proc}}(H_{0})$. Thus $\tilde{E}_{\tau}$ and $\sup_{t}\tilde{E}_{t}$ satisfy the Ville bounds of Theorem˜4.5.